OpenStudy (anonymous):

Find the fifth roots of 243(cos 240° + i sin 240°). i have this: r^5=243 Theta=240 degrees root:3 240/5+(0+360/3)=168 240/5+(1+360/3)=169 240/5+(2+360/3=170 making the roots: 3, 168, 169,170

OpenStudy (anonymous):

@jim_thompson5910 @satellite73

OpenStudy (anonymous):

divide the angle by 5

OpenStudy (anonymous):

one answer is \(3\left(\cos(48)+i\sin(48)\right)\)

OpenStudy (anonymous):

r^5=243 Theta=240 degrees root:3 240/5+(0+360/5)=168 240/5+(1+360/5)=169 240/5+(2+360/5=170????????? making the roots: 3, 168, 169,170

OpenStudy (anonymous):

i have no idea what this means 240/5+(0+360/3)=168

OpenStudy (anonymous):

i think you may be confused the roots are not real numbers, they are complex numbers your answers will all look like \(a+bi\) or \(r(\cos(\theta)+i\sin(\theta))\)

OpenStudy (anonymous):

and in each case when it looks like \[r(\cos(\theta)+i\sin(\theta))\] you will have \(r=3\) that doesn't change, only \(\theta\) changes

OpenStudy (anonymous):

there will be five answers because you are asked for the fifth roots the first angle you find by dividing \(240\) by \(5\) to get \(48\)

OpenStudy (anonymous):

wait now im super confused

OpenStudy (anonymous):

that is why i had the first answer above as \[3(\cos(48)+i\sin(48))\]

OpenStudy (anonymous):

lets go slow then this is the original question Find the fifth roots of \(243(\cos (240°) + i \sin (240°))\) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok is it clear to you that this is a complex number? a number that we could also write as \(a+bi\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in fact lets do it since \(\cos(240)=-\frac{1}{2}\) and \(\sin(240)=-\frac{\sqrt3}{2}\) this number is \[-\frac{232}{2}-\frac{232\sqrt3}{2}i\] not that this helps us find the answer, just so we understand that we are looking for five numbers \(a+bi\) with \((a+bi)^5=-\frac{232}{2}-\frac{232\sqrt3}{2}i\)

OpenStudy (anonymous):

at least we see now that \(3\) makes no sense as an answer, because \(3^5=243\) not \(-\frac{243}{2}-\frac{243\sqrt3}{2}i\)

OpenStudy (anonymous):

but to find the 5 answers in trig from is real easy just divide the angle by 5 the angle is \(240\) and \(240\div 5=48\) so one answer is \[3(\cos(48)+i\sin(48))\]

OpenStudy (anonymous):

and you are not even expected to write this number as \(a+bi\) because you do not know what \(\cos(48)\) is just leave it in trig form as it is

OpenStudy (anonymous):

that is one of the five answers to get the next one, add \(360\) to \(240\) get \(600\) then divide that by \(5\) and get \(120\)

OpenStudy (anonymous):

then the next complex number will be \[3(\cos(120)+i\sin(120))\]

OpenStudy (anonymous):

so i just basically keep doing that 5 times?

OpenStudy (anonymous):

yes keep adding 360 on the fifth time you will be where you started, at an angle co-terminal with 48

OpenStudy (anonymous):

for the next one do i add 360 to 600?

OpenStudy (anonymous):

@satellite73